# ABD: Default–Exception Abduction in Finite First-Order Worlds

### v1

@@ Abstract
Because domains are finite, validity and costs are computed via SMT (Z3), enabling exact verification and controlled difficulty. We evaluate ten frontier LLMs on 600 instances spanning all three scenarios and seven default theories. The best models achieve 90-99% training validity, but parsimony gaps of ${\sim}1$$1.5$ extra exceptions per world remain. Holdout evaluation reveals distinct generalization profiles: in ABD-Full and ABD-Partial, the dominant failure is parsimony inflation; in ABD-Skeptical, it is validity brittleness—rules that work on training often break on holdouts—while survivors show smaller gap inflation.
@@ 1 Introduction
We evaluate ten frontier models across all three regimes. Holdout evaluation reveals two distinct failure modes: parsimony inflation in ABD-Full/Partial (gaps roughly double on fresh worlds) and validity brittleness in ABD-Skeptical (rules that work on training break on holdouts, though survivors show smaller gap inflation). Models differ qualitatively in their performance profiles, pointing to different reasoning strategies. Details are in Section 6.
@@
  • We evaluate ten frontier models—Opus-4.6, Grok4, Grok4.1-fast, Gemini-3.1-Pro, Gemini-3-Pro, GPT-5.2, DeepSeek-Reasoner (DSR), Kimi-K2-Thinking, GPT-4o, and Hermes4222Model references: Anthropic Claude Opus 4.6 (?); xAI Grok 4 (?) and Grok 4.1-fast (?); Google Gemini 3 Pro (?) and Gemini 3.1 Pro (?); OpenAI GPT-5.2 (?); DeepSeek-Reasoner (?); Moonshot Kimi K2 Thinking (?); OpenAI GPT-4o (?); Nous Research Hermes 4 (?).—and provide analyses of validity, parsimony gaps, formula complexity, and holdout generalization. Gemini-3.1-Pro was released close to the conference deadline (on February 19, 2026). We chose to include it alongside Gemini-3-Pro: the recent model substantially improves validity while slightly losing in gap metrics.
@@ 6 Evaluation and Experiments
Table 3: Abduction Training Performance. Gap = normalized extra exceptions above solver lower bound (free-$\mathsf{Ab}$ baseline); GapG = vs gold reference (lower is better). Best values bold.
@@
Model V% Gap GapG V% Gap GapG V% Gap GapG V% Gap GapG
Opus-4.6 100 1.26 0.74 98 1.02 0.55 99 0.86 0.17 99 1.05 0.51
Grok4 88 0.92 0.41 91 1.54 1.06 89 1.08 0.40 89 1.22 0.68
GPT-5.2 85 0.93 0.42 93 1.30 0.82 100 1.49 0.80 92 1.24 0.70
Gemini-3.1 99 1.27 0.76 97 1.41 0.93 98 1.18 0.50 98 1.30 0.76
Gemini-3 77 1.24 0.72 66 1.22 0.77 93 1.25 0.59 77 1.23 0.70
Grok4.1f 94 1.35 0.84 95 1.70 1.23 96 1.46 0.82 95 1.53 1.00
DSR 94 1.45 0.93 96 1.54 1.07 93 1.73 1.05 95 1.56 1.02
Kimi-K2t 57 1.48 0.96 72 1.69 1.22 80 2.46 1.77 70 1.87 1.32
Hermes4 3 2.85 2.30 1 1.78 1.44 4 3.11 2.56 2 2.81 2.29
GPT-4o 9 3.05 2.59 23 3.19 2.68 29 2.87 2.32 20 3.04 2.52
@@
Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
Theory V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG
T1 100 0.81 92 0.95 97 1.10 97 1.03 97 0.99 97 1.46 95 1.46 74 2.04 3 4.00 16 5.29
T2 99 0.74 92 0.76 93 0.82 99 0.93 80 0.72 98 1.03 95 1.11 73 1.35 2 1.01 16 2.01
T3 99 0.71 90 0.74 88 0.66 98 0.71 71 0.65 95 1.05 97 0.98 69 1.42 1 2.00 28 1.25
T4 99 0.47 86 0.71 89 0.77 97 0.89 61 0.82 91 1.16 93 1.14 58 1.46 0 14 4.66
T5 98 0.16 90 0.60 97 0.61 98 0.56 92 0.54 93 0.81 97 0.87 80 1.00 9 2.75 5 5.17
T6 100 -0.88 86 -0.71 100 -0.88 100 -0.81 93 -0.79 100 -0.61 93 -0.65 64 -0.62 0 93 1.88
T7 96 0.03 92 0.19 100 0.24 100 0.24 96 1.07 96 0.64 88 0.63 84 1.12 4 1.00 60 1.67
@@
Gemini-3 76.8% 50.0% 1.23 2.13 +0.80 18.9
Grok4.1f 94.8% 81.3% 1.53 2.58 +0.97 11.1
DSR 94.7% 80.2% 1.56 2.66 +0.98 10.8
Kimi-K2t 69.5% 65.0% 1.87 2.82 +0.93 11.5
Hermes4 2.3% 10.5% 2.81 1.33 +1.19 9.0
GPT-4o 19.8% 20.8% 3.04 4.03 +0.99 8.6
@@ 6.1 Prompting and Inference Setup
All models are queried at temperature $T=0.1$ for near-deterministic outputs. For models offering a thinking mode, we used the provider-default reasoning configuration with 64K token budget; other models were queried in standard generation mode. Each of the 600 benchmark instances is presented independently (single-turn, no multi-round refinement). All results correspond to API snapshots from January–February 2026; model behavior may differ in future versions.
Each model outputs a candidate $\hat{\alpha}(x)$ in S-expression syntax. We compute training validity and exception cost as defined in Section 4, with cost semantics varying by scenario. All gap values reported are normalized by the number of worlds to keep scores comparable across instances.
@@ 6.2 Main Results
Table 3 summarizes training performance across models along two axes: (i) validity (does the repaired theory have any model?), and (ii) parsimony (how close is the exception count to a solver lower bound?). The table reports both Gap (vs. the relaxed free-$\mathsf{Ab}$ lower bound; see Section 4.1) and GapG (vs. planted gold rule). Two patterns stand out.
Several frontier models achieve high training validity (above 90%, with the best reaching nearly 99%), while the older baseline GPT-4o remains far behind. This indicates that ABD is not a “syntax-only” benchmark: producing a well-formed formula is easy, but producing a globally valid repair policy across multiple worlds is not.
Parsimony Separates Strong Models.
Among valid predictions, the dominant source of separation is cost. Using normalized gaps (extra exceptions per world), the strongest models cluster around roughly 1.0–1.6 additional abnormalities beyond the lower bound, whereas weaker baselines incur substantially larger gaps. Interpreting the normalized scale: with $\approx 6$1010 training worlds per instance (Table 2), a gap of $\approx 1.1$ corresponds to roughly 771111 extra abnormal elements per instance.
Several models “beat gold” (negative gold-relative gap) on a nontrivial fraction of instances (Table 11). Gold rules are generator anchors, not ground truth; beating gold often correlates with larger AST size. This signal is useful for refining the competitor pool and identifying gold rules that are overly complex, but it does not indicate “winning” the task—parsimony relative to the solver lower bound remains the primary metric.
@@ 6.3 Scenario and Theory Breakdown
A notable qualitative contrast in Table 3 is that some models benefit from ABD-Partial (higher validity), while others lose validity under partial observation. This suggests that the partial setting genuinely probes reasoning about what could be true rather than merely making the task easier.
@@
Across theories (Table 4), T5 is consistently easier, while T3/T4 exhibit larger spread across models. A plausible explanation is that theories with stronger relational structure (e.g., nested quantification or additional dependencies in consequents) constrain the exception rule more tightly, reducing the effectiveness of compact heuristics. This theory-conditioned variation is valuable: it yields a natural set of controlled “difficulty knobs” without changing the output language or evaluation protocol. Additionally, T6 exhibits negative GapG values for most frontier models (Table 4), indicating that model predictions frequently beat the planted gold rule; this suggests T6’s gold rules may be overly complex and flags these instances for generator refinement.
@@ 6.4 Holdout Generalization
Table 5 aggregates across all three scenarios and mixes two effects: (i) whether a rule stays satisfiable on holdouts (validity brittleness), and (ii) how parsimonious it remains when satisfiable (parsimony inflation). Table 6 disentangles these by reporting holdout validity conditioned on training validity. These two failure modes manifest differently across scenarios, as we show next.
@@
Gemini-3 461 461 290 62.9%
Grok4.1f 542 541 464 85.8%
DSR 568 567 480 84.7%
Kimi-K2t 417 415 379 91.3%
Hermes4 14 14 14 100.0%
@@
Opus-4.6 100% 87% 1.26 2.72 +1.41 98% 50% 1.02 2.12 +0.95 99% 49% 0.86 1.59 +0.25
Grok4 88% 61% 0.92 1.86 +0.96 91% 72% 1.54 2.79 +1.14 89% 38% 1.08 1.85 +0.43
GPT-5.2 85% 55% 0.93 2.73 +1.48 93% 67% 1.30 2.52 +1.03 100% 61% 1.49 2.06 -0.02
Gemini-3.1 99% 87% 1.27 2.60 +1.35 97% 68% 1.41 2.44 +1.00 98% 52% 1.18 1.80 +0.21
Gemini-3 77% 60% 1.24 2.39 +1.16 66% 41% 1.22 2.21 +0.98 93% 52% 1.25 1.66 +0.11
Grok4.1f 94% 86% 1.35 2.67 +1.31 95% 85% 1.70 2.85 +1.14 96% 69% 1.46 1.83 -0.02
DSR 94% 83% 1.45 2.92 +1.44 96% 85% 1.54 2.70 +1.10 93% 70% 1.73 2.23 +0.09
Kimi-K2t 57% 58% 1.48 2.85 +1.46 72% 67% 1.69 2.83 +1.16 80% 70% 2.46 2.78 +0.10
Hermes4 3% 5% 2.85 5.40 +1.42 1% 1% 1.78 3.00 +1.22 4% 31% 3.11 0.47 +0.96
GPT-4o 9% 17% 3.05 5.58 +2.03 23% 16% 3.19 3.31 +0.85 29% 31% 2.87 3.57 +0.83
@@
Theory Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
T1 +0.5 (19) +1.5 (11) +0.4 (19) +0.6 (17) +0.5 (16) +0.5 (23) +0.4 (24) +0.2 (26) +0.0 (1) +1.8 (1)
T2 +1.5 (108) +1.1 (102) +1.4 (103) +1.5 (126) +1.2 (80) +1.3 (132) +1.5 (127) +1.5 (105) +2.1 (3) +1.7 (22)
T3 +1.0 (86) +1.1 (81) +1.0 (65) +1.0 (87) +0.7 (54) +1.0 (102) +1.0 (93) +1.0 (74) +2.4 (1) +1.0 (33)
T4 +0.9 (81) +1.0 (81) +0.9 (93) +0.9 (96) +0.9 (60) +1.0 (114) +1.0 (132) +1.0 (84) +0.8 (4)
T5 +0.6 (61) +0.6 (61) +0.3 (67) +0.5 (73) +0.5 (64) +0.5 (67) +0.4 (80) +0.4 (67) +0.9 (8) +1.4 (1)
T6 -0.2 (3) -0.3 (2) +0.5 (4) -0.3 (2) +0.1 (4) +0.4 (7) +0.1 (5) +0.0 (2) +0.2 (12)
T7 +0.3 (13) +0.3 (12) +0.3 (16) +0.3 (16) +0.4 (12) +0.2 (19) +0.4 (19) +0.2 (21) +0.8 (1) +0.4 (15)

### v2

@@ Abstract
Because domains are finite, validity and costs are computed via SMT (Z3), enabling exact verification and controlled difficulty. We evaluate eleven frontier LLMs on 600 instances spanning all three scenarios and seven default theories. Among the strongest high-validity models, training cost gaps of ${\sim}1$$1.5$ extra exceptions per world remain and valid-formula AST stays in the low teens, while GPT-5.4 is a low-gap, high-AST outlier with lower validity. Holdout evaluation reveals distinct generalization profiles: in ABD-Full and ABD-Partial, the dominant failure is parsimony inflation; in ABD-Skeptical, it is validity brittleness—rules that work on training often break on holdouts—while survivors show smaller gap inflation.
@@ 1 Introduction
We evaluate eleven frontier models across all three regimes. Holdout evaluation reveals two distinct failure modes: parsimony inflation in ABD-Full/Partial (gaps roughly double on fresh worlds) and validity brittleness in ABD-Skeptical (rules that work on training break on holdouts, though survivors show smaller gap inflation). The models also separate into distinct performance profiles: Opus-4.6, Gemini-3.1, DSR, and Grok4.1f form a high-validity cluster; GPT-5.4 attains the best gap metrics but with very large formulas and weak holdout survival; Kimi-K2t is comparatively compact and conditionally robust but substantially less parsimonious. Details are in Section 6.
@@
  • We evaluate eleven frontier models—Opus-4.6, Grok4, Grok4.1-fast, Gemini-3.1-Pro, Gemini-3-Pro, GPT-5.2, GPT-5.4, DeepSeek-Reasoner (DSR), Kimi-K2-Thinking, GPT-4o, and Hermes4222Model references: Anthropic Claude Opus 4.6 (?); xAI Grok 4 (?) and Grok 4.1-fast (?); Google Gemini 3 Pro (?) and Gemini 3.1 Pro (?); OpenAI GPT-5.2 (?) and GPT-5.4 (?); DeepSeek-Reasoner (?); Moonshot Kimi K2 Thinking (?); OpenAI GPT-4o (?); Nous Research Hermes 4 (?).—and provide analyses of validity, parsimony gaps, formula complexity, and holdout generalization. Gemini-3.1-Pro was released close to the conference deadline (on February 19, 2026). We chose to include it alongside Gemini-3-Pro: the newer model substantially improves validity (98.0% vs. 76.8% on train; 69.7% vs. 50.0% on holdout) with only a modest increase in train gap (1.30 vs. 1.23).
@@ 6 Evaluation and Experiments
Table 3: Abduction Training Performance. Models are grouped by overall repaired train validity into high ($90\%$), intermediate, and low ($50\%$) bands; within each band, rows are ordered by overall Gap. V% = train-valid after conservative suffix repair when applicable; SV% = strict train-validity without any repair; AST = mean abstract-syntax-tree size among train-valid formulas; Gap = normalized extra exceptions above the solver lower bound (free-$\mathsf{Ab}$ baseline); GapG = normalized extra exceptions vs. the planted gold reference. Lower is better for AST, Gap, and GapG. Best values bold.
@@
Model V% SV% AST Gap GapG V% SV% AST Gap GapG V% SV% AST Gap GapG V% SV% AST Gap GapG
High Validity ($90\%$)
Opus-4.6 100 100 12.4 1.26 0.74 98 98 15.9 1.02 0.55 99 99 15.8 0.86 0.17 99 99 14.7 1.05 0.51
GPT-5.2 85 84 36.6 0.93 0.41 92 92 16.3 1.30 0.84 100 100 15.7 1.49 0.80 92 92 22.0 1.25 0.70
Gemini-3.1 99 99 12.5 1.27 0.76 97 97 11.4 1.40 0.93 98 98 13.5 1.18 0.50 98 98 12.3 1.30 0.76
Grok4.1f 96 94 12.8 1.32 0.80 95 95 9.0 1.69 1.23 96 96 11.8 1.46 0.82 95 95 11.0 1.51 0.98
DSR 96 94 11.9 1.42 0.90 96 96 9.4 1.54 1.08 96 93 11.7 1.69 1.01 96 95 10.8 1.55 1.00
Intermediate Validity
GPT-5.4 76 68 106.1 0.13 -0.38 86 81 61.5 0.65 0.18 94 92 32.6 0.37 -0.28 85 80 65.9 0.42 -0.12
Grok4 87 87 26.2 0.90 0.39 92 92 10.5 1.55 1.09 89 89 16.8 1.08 0.40 89 89 17.1 1.21 0.67
Gemini-3 77 77 18.1 1.23 0.72 65 65 16.5 1.22 0.79 93 93 15.3 1.25 0.59 77 77 16.6 1.23 0.70
Kimi-K2t 60 57 13.3 1.45 0.93 73 72 10.3 1.68 1.21 83 80 10.3 2.41 1.74 72 70 11.1 1.84 1.30
Low Validity ($50\%$)
Hermes4 3 3 8.2 2.85 2.31 1 1 8.0 1.78 1.44 4 4 6.5 3.11 2.56 2 2 7.4 2.80 2.29
GPT-4o 9 9 8.4 3.05 2.59 23 23 8.4 3.18 2.68 29 29 4.7 2.87 2.32 20 20 6.9 3.04 2.52
@@
Opus-4.6 GPT-5.2 Gemini-3.1 Grok4.1f DSR GPT-5.4 Grok4 Gemini-3 Kimi-K2t Hermes4 GPT-4o
Theory V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG
T1 100 0.81 97 1.10 97 1.03 97 1.46 95 1.46 95 0.07 92 0.95 97 0.99 76 1.97 3 4.00 16 5.29
T2 99 0.74 94 0.84 99 0.93 98 1.05 96 1.10 82 0.05 91 0.76 81 0.71 75 1.33 2 1.01 16 2.00
T3 99 0.71 88 0.67 98 0.71 97 1.02 97 0.97 82 -0.14 90 0.74 70 0.63 74 1.36 1 2.00 28 1.25
T4 99 0.46 88 0.78 97 0.89 91 1.12 95 1.13 86 -0.18 86 0.71 60 0.86 59 1.44 0 14 4.66
T5 98 0.16 96 0.61 98 0.56 92 0.80 99 0.86 85 -0.21 90 0.62 92 0.53 82 0.99 8 2.82 5 5.17
T6 100 -0.88 100 -0.88 100 -0.81 100 -0.61 93 -0.65 93 -0.99 86 -0.71 93 -0.79 64 -0.62 0 93 1.88
T7 96 0.03 100 0.24 100 0.24 96 0.64 92 0.60 96 -0.05 92 0.19 96 1.07 88 1.16 4 1.00 60 1.67
@@
Grok4.1f 95.2% 82.2% 1.51 2.54 +0.94 11.3
DSR 95.9% 80.6% 1.55 2.64 +0.96 11.1
GPT-5.4 85.2% 24.8% 0.42 1.37 +0.56 73.1
Grok4 89.4% 59.1% 1.21 2.32 +0.95 17.6
Gemini-3 76.9% 49.6% 1.23 2.13 +0.80 18.8
Kimi-K2t 71.5% 64.8% 1.84 2.85 +0.93 12.3
Hermes4 2.3% 3.0% 2.80 4.71 +1.17 8.8
GPT-4o 19.8% 19.0% 3.04 4.01 +0.99 8.6
@@ 6.1 Prompting and Inference Setup
All models are queried at temperature $T=0.1$ for near-deterministic outputs. For models offering a thinking mode, we used the provider-default reasoning configuration with 64K token budget; other models were queried in standard generation mode. Each of the 600 benchmark instances is presented independently (single-turn, no multi-round refinement). All results correspond to API snapshots from January–March 2026 (through March 7 for GPT-5.4); model behavior may differ in future versions.
Each model outputs a candidate $\hat{\alpha}(x)$ in S-expression syntax. We compute training validity and exception cost as defined in Section 4, with cost semantics varying by scenario. All gap values reported are normalized by the number of worlds to keep scores comparable across instances. When the only parse failure is a suffix of missing right parentheses, we conservatively auto-close the formula at end of string, evaluate the repaired form, and report such cases separately from unrepaired parse errors.
@@ 6.2 Main Results
Table 3 summarizes training performance across models along three axes: (i) validity (does the repaired theory have any model?), and (ii) syntactic compactness (how large are the train-valid formulas?), and (iii) cost parsimony (how close is the exception count to a solver lower bound?). The table reports repaired validity (V%), strict non-repaired validity (SV%), AST, Gap (vs. the relaxed free-$\mathsf{Ab}$ lower bound; see Section 4.1), and GapG (vs. planted gold rule). Two patterns stand out.
Frontier models already separate into distinct performance profiles. Opus-4.6, Gemini-3.1, DSR, and Grok4.1f combine high training validity with relatively compact valid formulas, while GPT-5.4 pushes Gap and GapG much lower only by accepting substantially lower validity and much larger formulas. This contrast matters because formula size is not merely cosmetic: the largest-AST systems generalize much worse on holdout, suggesting that some models reduce train cost through brittle case-splitting rather than portable repair rules.
Cost Gap and Formula Size Separate Different Models.
Among valid predictions, semantic cost and syntactic compactness are only partially aligned. Among the strongest high-validity models, normalized gaps cluster around roughly 1.0–1.6 additional abnormalities beyond the lower bound, while mean AST among valid formulas stays in the low teens (10.8–14.7 for Opus-4.6, Gemini-3.1, DSR, and Grok4.1f). GPT-5.4 is the main exception: it attains the best overall train Gap (0.42) and GapG (-0.12), but this cost advantage comes with lower validity (85.2%) and by far the largest valid formulas (AST 65.9 overall). Low cost gap is therefore not by itself a sufficient notion of parsimony. Interpreting the normalized scale: with $\approx 6$1010 training worlds per instance (Table 2), a gap of $\approx 1.1$ corresponds to roughly 771111 extra abnormal elements per instance.
Several models “beat gold” (negative gold-relative gap) on a nontrivial fraction of instances (Table 11). Gold rules are generator anchors, not ground truth; beating gold often correlates with larger AST size. GPT-5.4 makes this especially clear: it beats gold on 56.8% of valid FULL/PARTIAL instances, but those gold-beating formulas have mean AST 90.1. This signal is useful for refining the competitor pool, but it does not indicate “winning” the task—cost parsimony relative to the solver lower bound remains the primary semantic metric, while AST tracks whether that cost is achieved with compact formulas.
@@ 6.3 Scenario and Theory Breakdown
A notable qualitative contrast in Table 3 is that some models benefit from ABD-Partial (higher validity), while others lose validity under partial observation. This suggests that the partial setting genuinely probes reasoning about what could be true rather than merely making the task easier. Gemini-3.1 improves sharply over Gemini-3 in every regime, especially on validity. DSR and Grok4.1f combine high validity with low-teens valid-formula AST and later emerge as the strongest holdout generalizers. Kimi-K2t shows the opposite tradeoff from GPT-5.4: once it finds a train-valid rule, it is unusually robust on holdout, but its gaps remain large. GPT-5.4 drives cost down with much larger repair policies, which rarely survive holdout unchanged.
@@
Across theories (Table 4), T5 is consistently easier, while T3/T4 exhibit larger spread across models. A plausible explanation is that theories with stronger relational structure (e.g., nested quantification or additional dependencies in consequents) constrain the exception rule more tightly, reducing the effectiveness of compact heuristics. This theory-conditioned variation is valuable: it yields a natural set of controlled “difficulty knobs” without changing the output language or evaluation protocol. GPT-5.4 also attains the best GapG in every theory row without leading validity on any theory family, reinforcing that low train cost and robust validity are distinct capabilities. Additionally, T6 exhibits negative GapG values for most frontier models (Table 4), indicating that model predictions frequently beat the planted gold rule; this suggests T6’s gold rules may be overly complex and flags these instances for generator refinement.
@@ 6.4 Holdout Generalization
[…]
@@
Grok4.1f 512 511 440 86.1%
DSR 543 542 455 83.9%
GPT-5.4 511 509 145 28.5%
Grok4 505 504 329 65.3%
Gemini-3 435 435 272 62.5%
Kimi-K2t 429 427 383 89.7%
Hermes4 13 13 13 100.0%
@@
Opus-4.6 100% 87% 1.26 2.66 +1.34 98% 51% 1.02 2.09 +0.92 99% 49% 0.86 1.59 +0.25
GPT-5.2 85% 54% 0.93 2.74 +1.49 92% 68% 1.30 2.52 +1.02 100% 61% 1.49 2.06 -0.02
Gemini-3.1 99% 87% 1.27 2.60 +1.35 97% 68% 1.40 2.44 +1.00 98% 52% 1.18 1.80 +0.21
Grok4.1f 96% 88% 1.32 2.59 +1.27 95% 85% 1.69 2.85 +1.14 96% 69% 1.46 1.83 -0.02
DSR 96% 85% 1.42 2.86 +1.42 96% 85% 1.54 2.72 +1.11 96% 70% 1.69 2.23 +0.09
GPT-5.4 76% 18% 0.13 1.07 +0.68 86% 25% 0.65 2.13 +0.96 94% 33% 0.37 0.73 +0.04
Grok4 87% 61% 0.90 1.86 +0.98 92% 72% 1.55 2.80 +1.12 89% 38% 1.08 1.85 +0.43
Gemini-3 77% 60% 1.23 2.42 +1.21 65% 40% 1.22 2.21 +0.98 93% 52% 1.25 1.66 +0.11
Kimi-K2t 60% 58% 1.45 2.89 +1.45 73% 66% 1.68 2.87 +1.16 83% 71% 2.41 2.76 +0.10
Hermes4 3% 5% 2.85 5.51 +1.40 1% 1% 1.78 3.00 +1.22 4% 4% 3.11 4.07 +0.96
GPT-4o 9% 17% 3.05 5.58 +2.03 23% 16% 3.18 3.31 +0.86 29% 25% 2.87 3.38 +0.83
@@
Theory Opus-4.6 GPT-5.2 Gemini-3.1 Grok4.1f DSR GPT-5.4 Grok4 Gemini-3 Kimi-K2t Hermes4 GPT-4o
T1 +0.5 (19) +0.4 (19) +0.6 (17) +0.5 (23) +0.4 (24) +0.6 (7) +1.5 (11) +0.5 (16) +0.2 (26) +0.0 (1) +1.8 (1)
T2 +1.5 (109) +1.3 (97) +1.5 (126) +1.3 (123) +1.5 (119) +0.8 (45) +1.1 (94) +1.2 (76) +1.5 (105) +2.1 (3) +1.7 (22)
T3 +1.0 (86) +1.0 (61) +1.0 (87) +1.0 (99) +1.1 (88) +0.8 (24) +1.1 (76) +0.7 (50) +1.0 (76) +2.4 (1) +1.0 (33)
T4 +0.8 (82) +0.8 (83) +0.9 (96) +0.9 (105) +0.9 (122) +0.4 (30) +1.0 (75) +0.9 (53) +1.0 (85) +0.8 (4)
T5 +0.6 (61) +0.3 (64) +0.5 (73) +0.5 (64) +0.4 (78) +0.2 (27) +0.6 (59) +0.5 (61) +0.5 (68) +0.8 (7) +1.4 (1)
T6 -0.2 (3) +0.5 (4) -0.3 (2) +0.4 (7) +0.1 (5) +0.3 (3) -0.3 (2) +0.1 (4) +0.0 (2) +0.2 (12)
T7 +0.3 (13) +0.3 (16) +0.3 (16) +0.2 (19) +0.4 (19) +0.1 (9) +0.3 (12) +0.4 (12) +0.2 (21) +0.8 (1) +0.4 (15)